OpenStudy (j.f):

The first three terms of a progression are 3, 6, 12,.. what is the value of the fourth term. I know how to find the value but would the question be set up as 3-6-12=x?

OpenStudy (j.f):

24 a=a+(4-1)d= 6+3(6)= 24

OpenStudy (mathmath333):

yes \(24\) is correct \(3(2^0+2^1+2^2+2^3\ldots\)